The decomposition of SO2Cl2 is first order in SO2Cl2 and has a rate constant of 1.36×10−4 s−1 at a certain temperature.

Part A:

If the initial concentration of SO2Cl2 is 1.00 M, how long will it take for the concentration to decrease to 0.78 M ?

Part B:

If the initial concentration of SO2Cl2 is 0.125 M , what is the concentration of SO2Cl2 after 180 s?

Part C:

If the initial concentration of SO2Cl2 is 0.125 M , what is the concentration of SO2Cl2 after 520 s?
Answer #1

part A

k = 1/tln(a/x)

1.36*10^-4 = (1/t)ln(1/0.78)

t = 1826.92 sec

part B

1.36*10^-4 = (1/180)ln(0.125/x)

x = 0.122 M

part C

1.36*10^-4 = (1/520)ln(0.125/x)

x= 0.116 M